# Latent Inter-User Difference Modeling for LLM Personalization

### v1

@@ Abstract
Yilun Qiu1, Tianhao Shi2, Xiaoyan Zhao3, Fengbin Zhu1, Yang Zhang1, Fuli Feng2 1National University of Singapore 2University of Science and Technology of China 3The Chinese University of Hong Kong qiuyilun@u.nus.edu, sth@mail.ustc.edu.cn, xzhao@se.cuhk.edu.hk, zhfengbin@gmail.com, zyang1580@gmail.com, fulifeng93@gmail.com
@@ 1 Introduction
With continuous advancements in general-purpose intelligence, large language models (LLMs) Achiam et al. (2023); Grattafiori et al. (2024); Yang et al. (2025); Team et al. (2025); Guo et al. (2025) are increasingly integrated into everyday life, assisting users in making decisions Yao et al. (2023); Zhao et al. (2025c), retrieving information Zhao et al. (2024a); Asai et al. (2024), and task management Shen et al. (2024b). This growing presence has raised expectations for LLMs to go beyond generic, one-size-fits-all responses and instead produce responses that align with individual users’ unique preferences. To meet these heightened expectations, there has been a growing interest in LLM personalization  Chen et al. (2024b); Zhang et al. (2024b); Xu et al. (2025); Liu et al. (2025b), which aims at tailoring model outputs based on user-specific information.
@@
  • We introduce a novel method, DEP, to achieve latent difference modeling, equipped with a sparse autoencoder to extract task-relevant differences while filtering out noise.
@@ 2 Preliminary
where $\hat{y}_{u^{\prime}}^{i^{\prime}}$ denotes the generated text, $D_{u^{\prime}}$ denotes the history of the target user $u^{\prime}$, and $\phi({D_{u^{\prime}};D})$ denotes the process that extracts user-specific and difference-aware preference signals from $D_{u}$ and $D$. This memory retrieval framework supports lifelong user modeling without requiring LLM retraining Zhuang et al. (2024), making it both adaptable and resource-efficient for real-world personalization scenarios.
@@ 3.2 Difference-aware Embedding-based Personalization (DEP)
In this section, we introduce three key steps of DEP: constructing difference-aware representations, distilling them via a sparse autoencoder, and injecting them into an LLM for personalization.
@@ 3.2.1 Latent-space Difference-aware Representation Modeling
where $e_{his}^{i}$ denotes the user-specific embedding that reflects the preference pattern of user $u^{\prime}$ on item $y$. Next, to construct inter-user embeddings, we identify the set of peer users who also interacted with item $i$, excluding $u^{\prime}$, as $\{u_{1},u_{2},\dots,u_{m}\}$, where $u_{j}$ denotes the $j$-th peer user of item $i$. Each peer user $u_{j}$ provides a review $y^{i}_{u_{j}}$, which is encoded into an embedding:
@@ 3.2.3 Representation Injection
[…]
@@ 4.2 Main Results (RQ1)
  • Incorporating user context significantly improves the model’s capability for personalized text generation. Methods like RAG and PAG leverage retrieved user information to condition generation, significantly outperforming the Non-Perso baseline. DPL further improves upon these by explicitly modeling inter-user differences, achieving the relatively best performance among all ICL-based methods. This shows that capturing user differences yields better personalization than simple relevance or summarization.
@@ 4.4.1 Impact of History Number (RQ3)
For a more comprehensive view, we provide the full results across other evaluation metrics and datasets in Appendix D.3.
@@ 4.4.2 Impact of User Uniqueness (RQ4)
Following the procedure in DPL Qiu et al. (2025), we further investigate how user uniqueness affects personalization performance. Similarly, we adopt a grouping strategy based on the user embedding derived from historical reviews. Specifically, we compute the Euclidean distance between each user’s review embedding and the global average embedding across all users, and divide users into two groups: the top 50% as Unique users and the bottom 50% as Non-Unique users.
@@ 5 Related Work
Recent advancements Brown et al. (2020); Shen et al. (2024a); Liang et al. ; Achiam et al. (2023); Zhang et al. (2023); Zhao et al. (2024b); Zhang et al. (2024a) in large language models (LLMs) have demonstrated their strong generalization capabilities across diverse tasks Zheng et al. (2023); Zhang et al. (2025b, c); Du et al. (2025); Liu et al. (2025d); Chen et al. (2025b); Fang et al. (2025); Liu et al. (2025a). However, their ability to reflect personalized user intent remains limited. Consequently, the personalization of LLMs has become a critical research direction, aiming to adapt general-purpose models to individual user preferences Kirk et al. (2024); Chen et al. (2025a); Lin et al. (2025); Zhao et al. (2025a, b); Mok et al. (2025); Zhao et al. (2025d). Among various approaches, the memory-retrieval framework Salemi et al. (2024) is widely adopted for its interpretability and scalability. It retrieves user-specific signals from interaction history to guide the model without changing its parameters. Methods under this framework generally fall into two types: retrieval-augmented generation (RAG) and profile-augmented generation (PAG). RAG-based approaches retrieve relevant past interactions to construct a personalized prompt. For example, HYDRA Zhuang et al. (2024) employs a personalized reranker to refine retrieval quality, while PERAL Mysore et al. (2024) trains a retriever with a scale-calibrated objective to select useful information. In contrast, PAG-based methods summarize the user’s behavior into a condensed profile, which is then integrated into the prompt to guide generation Richardson et al. (2023b).
@@ 6 Conclusion
While our proposed method DEP demonstrates strong performance in personalized text generation, it also introduces several limitations. First, the method relies on sufficient user history to construct meaningful embeddings; in cold-start or data-sparse settings, its effectiveness may degrade. Second, although more efficient than language-based comparison methods, the computation of difference-aware embeddings and sparse autoencoding introduces additional overhead compared to standard prompting pipelines. Lastly, our evaluation is centered on review generation, where preferences are explicit; adapting the approach to broader tasks like dialogue or recommendation requires further study.
@@ Appendix B Baseline Details
    Training Dataset 3996 37.47$\pm$33.53 1608.82$\pm$1476.99
    Validation Dataset 512 39.14$\pm$36.01 1557.29$\pm$1378.43
    Test Dataset Books 317 34.84$\pm$22.55 1194.90$\pm$802.44
    Movies & TV 1925 41.11$\pm$35.90 1704.61$\pm$1752.44
    CDs & Vinyl 1754 38.50$\pm$32.37 1600.04$\pm$1419.89
  • PPlug Liu et al. (2025c): A plug-and-play personalization method that encodes a user’s history into a dense user-specific embedding through a lightweight user embedder. This embedding is constructed via input-aware attention over user histories. The resulting embedding, along with an instruction embedding, are projected into the LLM input space via a trainable projector and prepended to the input to guide a frozen LLM. In our implementation of PPlug, we adopt the same user embedder as used in our proposed method.
@@ C.2.1 Method Parameters
A maximum of 8 user history entries are retrieved for each instance. If the input exceeds the context length limit, excess histories are discarded to ensure compatibility.
@@ C.2.3 Inference Settings
Appendix D Complete Ablation Studies & In-Depth Analysis
@@ D.3 Impact of History Number
Appendix E Additional Experiment & Analysis
E.1 Retrieval Method
@@
E.2 System Prompt Guidance
As described in Section 6, we incorporate additional information into the system prompt to help the model better understand the injected personalization prompts. To assess its effectiveness, we conduct experiments to analyze the impact of this guidance. Table 9 reports the results across all datasets and evaluation metrics. We observe that incorporating system prompt guidance consistently improves performance across the board. Hence, we adopt the system prompt guidance by default in all experiments.
@@ Appendix F Overview of Templates & Prompts
In this section, we illustrate the prompt design used in our framework. As shown in Figure 6, the upper part depicts the system prompt, which defines the model’s global behavior and task instruction. The lower part shows an example of the input prompt, including retrieved user histories and object descriptions, which are fed into the model for generation. This prompt structure follows the retrieve-and-inject paradigm, where both user-specific and difference-aware embeddings are embedded via soft prompts [HIS_TOKEN_i] and [DIFF_TOKEN_i] to guide the generation. The four special tokens <his_token_start>, <his_token_end>, <diff_token_start>, and <diff_token_end> are introduced to explicitly mark the boundaries of user-specific and difference-aware embeddings in the input sequence.

### v2

@@ Abstract
Yilun Qiu1, Tianhao Shi2, Xiaoyan Zhao311footnotemark: 1, Fengbin Zhu1, Yang Zhang1thanks: Corresponding Authors, Fuli Feng2 1National University of Singapore 2University of Science and Technology of China 3The Chinese University of Hong Kong qiuyilun@u.nus.edu, sth@mail.ustc.edu.cn, xzhao@se.cuhk.edu.hk, zhfengbin@gmail.com, zyang1580@gmail.com, fulifeng93@gmail.com
@@ 1 Introduction
With continuous advancements in general-purpose intelligence, large language models (LLMs) Achiam et al. (2023); Grattafiori et al. (2024); Yang et al. (2025); Team et al. (2025); Guo et al. (2025) are increasingly integrated into everyday life, assisting users in making decisions Yao et al. (2023); Zhao et al. (2025d), retrieving information Zhao et al. (2024a); Fang et al. (2025b), and task management Shen et al. (2024a, c). This growing presence has raised expectations for LLMs to go beyond generic, one-size-fits-all responses and instead produce responses that align with individual users’ unique preferences. To meet these heightened expectations, there has been a growing interest in LLM personalization  Kirk et al. (2024); Chen et al. (2024b); Zhang et al. (2024b); Xu et al. (2025b); Liu et al. (2025b), which aims at tailoring model outputs based on user-specific information.
@@
  • We introduce a novel method, DEP, to achieve latent inter-user difference modeling, equipped with a sparse autoencoder to extract task-relevant differences while filtering out noise.
@@ 2 Preliminary
where $\hat{y}_{u^{\prime}}^{i^{\prime}}$ denotes the generated text, $D_{u^{\prime}}$ denotes the history of the target user $u^{\prime}$, and $\phi({D_{u^{\prime}};D})$ denotes the process that extracts user-specific and difference-aware preference signals from $D_{u}$ and $D$. This memory retrieval framework supports lifelong user modeling without requiring LLM retraining, making it both adaptable and resource-efficient for real-world personalization scenarios.
@@ 3.2 Difference-aware Embedding-based Personalization (DEP)
In this section, we introduce three key steps of DEP: constructing latent difference-aware representations, distilling them via a sparse autoencoder, and injecting them into an LLM for personalization.
@@ 3.2.1 Latent-space Difference-aware Representation Modeling
where $e_{his}^{i}$ denotes the user-specific embedding that reflects the preference pattern of user $u^{\prime}$ on item $i$. Next, to construct inter-user embeddings, we identify the set of peer users who also interacted with item $i$, excluding $u^{\prime}$, as $\{u_{1},u_{2},\dots,u_{m}\}$, where $u_{j}$ denotes the $j$-th peer user of item $i$. Each peer user $u_{j}$ provides a review $y^{i}_{u_{j}}$, which is encoded into an embedding:
@@ 3.2.3 Representation Injection
where $I^{*}_{u^{\prime}}$ denotes the top-$N$ retrieved items from the target user’s interacted history, and $\mathcal{S}(i^{\prime},\{i,p^{i}_{\text{his}},p^{i}_{\text{diff}}\}_{{i}\in{I^{*}_{u^{\prime}}}})$ is a textual prompt constructed from both the target item $i^{\prime}$ and the soft prompts to model inter-user differences, and the original user’s original review history to model user’s own writing patterns. The template can be found in Figure 6 in Appendix F.
@@ 4.2 Main Results (RQ1)
  • Incorporating context information significantly improves the model’s capability for personalized text generation. Methods like RAG and PAG leverage retrieved user information for generation, significantly outperforming the Non-Perso baseline. DPL further improves upon these by explicitly modeling inter-user differences, achieving the relatively best performance among all ICL-based methods. This shows that capturing user differences yields better personalization than simple relevance or summarization.
@@ 4.4.1 Impact of History Number (RQ3)
For a more comprehensive view, we provide more detailed results across other evaluation metrics and datasets in Appendix D.3.
@@ 4.4.2 Impact of User Uniqueness (RQ4)
Following the procedure in DPL, we further investigate how user uniqueness affects personalization performance. Similarly, we adopt a grouping strategy based on the user embedding derived from historical reviews. Specifically, we compute the Euclidean distance between each user’s review embedding and the global average embedding across all users, and divide users into two groups: the top 50% as Unique users and the bottom 50% as Non-Unique users.
@@ 5 Related Work
Recent advancements Zeng et al. (2025); Liang et al. ; Zhang et al. (2023, 2024a); Zhao et al. (2024b); Liu et al. (2025d); Chen et al. (2025b); Yao et al. (2025) in large language models (LLMs) have demonstrated their strong generalization capabilities across diverse tasks Zheng et al. (2023); Zhang et al. (2025b, c); Du et al. (2025); Liu et al. (2025a); Zhao et al. (2025b); He et al. (2025); Fang et al. (2025c, a); Sheng et al. (2025). However, their ability to reflect personalized user intent remains limited. Consequently, the personalization of LLMs has become a critical research direction, aiming to adapt general-purpose models to individual user preferences Kirk et al. (2024); Chen et al. (2025a); Lin et al. (2025); Mok et al. (2025); Zhao et al. (2025a, c, e); Shen et al. (2024b); Xu et al. (2025a, c). Among various approaches, the memory-retrieval framework Salemi et al. (2024) is widely adopted for its interpretability and scalability. It retrieves user-specific signals from interaction history to guide the model without changing its parameters. Methods under this framework generally fall into two types: retrieval-augmented generation (RAG) and profile-augmented generation (PAG). RAG-based approaches retrieve relevant past interactions to construct a personalized prompt. For example, HYDRA Zhuang et al. (2024) employs a personalized reranker to refine retrieval quality, while PERAL Mysore et al. (2024) trains a retriever with a scale-calibrated objective to select useful information. In contrast, PAG-based methods summarize the user’s behavior into a condensed profile, which is then integrated into the prompt to guide generation Richardson et al. (2023b).
@@ 6 Conclusion
While our proposed method DEP demonstrates strong performance in personalized text generation, it also introduces several limitations. First, the method relies on sufficient user history to construct meaningful embeddings; in cold-start or data-sparse settings, its effectiveness may degrade. Second, although more efficient than language-based comparison methods, the computation of difference-aware embeddings and the sparse autoencoder introduces additional overhead compared to standard prompting pipelines. Lastly, our evaluation is centered on review generation, where preferences are explicit; adapting the approach to broader tasks like dialogue or recommendation requires further study.
@@
We use AI assistants (e.g., ChatGPT) as auxiliary tools for writing refinement and coding support, while all research ideas, experimental designs, and final decisions are made by the authors.
@@ Appendix B Baseline Details
[…]
@@
  • PPlug Liu et al. (2025c): A plug-and-play personalization method that encodes a user’s history into a dense user-specific embedding through a lightweight user embedder. This embedding is constructed via input-aware attention over user histories. The resulting embedding, along with an instruction embedding, is projected into the LLM input space via a trainable projector and prepended to the input to guide a frozen LLM. In our implementation of PPlug, we adopt the same user embedder as used in our proposed method.
@@ C.2.1 Method Parameters
At most 8 user history entries are retrieved for each instance. If the input exceeds the context length limit, excess histories are discarded to ensure compatibility.
@@ C.2.3 Inference Settings
Appendix D Complete Ablation Studies & Additional Experiments
@@ D.3 Impact of History Number
D.4 Retrieval Method
@@
D.5 System Prompt Guidance
As shown in Figure 6, we incorporate additional information into the system prompt to help the model better understand the injected personalization prompts. To assess its effectiveness, we conduct experiments to analyze the impact of this guidance. Table 9 reports the results across all datasets and evaluation metrics. We observe that incorporating system prompt guidance consistently improves performance across the board. Hence, we adopt the system prompt guidance by default in all experiments.
Appendix E Further In-Depth Analysis
E.1 Interpretability Analysis
To further examine how personalization is achieved, we conducted an interpretability analysis comparing DEP with the baseline DPL. The results show that DEP better captures users’ word choice, semantic information, and overall writing patterns, thereby aligning generated reviews more closely with users’ authentic writing style.
Table 10 presents an illustrative example. Compared with DPL, which produces a more formal and detached review, DEP generates text that reflects the user’s actual phrasing, tone, and evaluative stance. This demonstrates DEP’s ability to internalize and reproduce the personalized linguistic patterns of users.
Source Review Text
DPL-generated Review On the positive side, the film does a good job of maintaining a level of respect for its audience by not overloading with graphic gore, which is a relief for those who might find that kind of content too disturbing.
DEP-generated Review I’ll give it that, it is not full of gore and disgust. If you are looking for a good horror film, this is not it.
User’s Actual Review Not something I’m eager to watch again, but in a pinch, it is at least not full of gore and disgust. I’ll give it that.
Table 10: Comparison between reviews generated by DPL and DEP and the user’s actual review.
E.2 Practical Applicability
To assess the practical applicability of our method, we further examine its training cost. Importantly, our approach does not involve tuning the LLM itself; instead, the backbone model remains fixed while only the input component is tuned. This design substantially reduces the number of trainable parameters to approximately 0.4% of the LLM’s total parameters. Consequently, the additional computational overhead is minimal. In practice, the training cost is comparable to the soft prompt tuning baseline PPlug, with both methods requiring around 40 minutes per epoch on our datasets using a single GPU.
@@ Appendix F Overview of Templates & Prompts
In this section, we illustrate the prompt design used in our framework. As shown in Figure 6, the upper part depicts the system prompt, which defines the model’s global behavior and task instruction. The lower part shows an example of the input prompt, including retrieved user histories and object descriptions, which are fed into the model for generation. This prompt structure follows the retrieve-and-inject paradigm, where both user-specific and difference-aware embeddings are embedded via soft prompts [HIS_TOKEN_i] and [DIFF_TOKEN_i] to guide the generation. The four special tokens <his_token_start>, <his_token_end>, <diff_token_start>, and <diff_token_end> are introduced to explicitly mark the boundaries of user-specific and difference-aware embeddings in the input sequence. The note part is the system prompt guidance described in Section D.5.